# Vectorized Computation of Euler Characteristic Functions and Transforms

### v2

@@ Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are either not optimized for speed or do not scale to higher-dimensional settings. In this work, we present a vectorized framework for computing such topological descriptors using tensor operations, which is highly optimized for GPU architectures and works in full generality across simplicial and cubical complexes of arbitrary dimension. Experimentally, the framework demonstrates significant speedups over existing methods when computing the WECT and ECF across a variety of two- and three-dimensional datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT.
@@ 1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant: the Euler characteristic [10]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [9] and Euler characteristic functions [20].
Topological transforms [23] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a directional filtration of $K$ is constructed and then a topological invariant is applied to that filtration. Applying the Euler characteristic (EC) to a single directional filtration results in the Euler characteristic function (ECF), while applying it to each directional filtration in $\mathbb{S}^{n-1}$ results in the Euler characteristic transform (ECT) [23, 15]. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [6, 11]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
Useful variants of the ECF and ECT are the weighted Euler characteristic function (WECF) and the weighted Euler characteristic transform (WECT) for weighted complexes [12]. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [12]. Moreover, the WECT has been employed in a broad range of applications, including: in oncology for predicting survival rates of patients with glioblastoma multiforme brain tumors [5], in plant biology for classifying seed phenotypes [1], and in image classification tasks [3, 12]. Due to its popularity, numerous software packages already exist for computing the WECT and its variants, including those presented in [1, 5, 13, 14, 17, 18]. While many of these packages provide useful tools, they are either not optimized for speed or are limited in the types (or dimension) of complexes they can process.
In this work, we provide a fully vectorized framework for computing WECFs and WECTs. Our implementation applies in full generality to arbitrary geometric simplicial and cubical complexes in Euclidean spaces. The vectorized operations are highly optimized for fast computation on modern GPU hardware, providing speedups over existing implementations.
@@ 2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to simplicial complexes in this paper; however, our results and code work for cubical complexes.
We use abstract simplicial complexes throughout this work. For a (abstract) simplicial complex $K$ and for any $i\in\{0,1,\ldots,\dim(K)\}$, we denote the set of $i$-simplices of $K$ by $K_{i}$ and treat $\sigma\in K$ as a set of vertices. Furthermore, we define $k=|K|$ and $k_{i}=|K_{i}|$.
2.1 Lower-Star Filtrations
Here, we provide the necessary background on filtrations.
Definition 1 (Filtrations and Filters).
Let $K$ be a simplicial complex and let $\operatorname{sub}(K)$ denote the set of subcomplexes of $K$. A filtration of $K$ is a function $F\colon\mathbb{R}\to\operatorname{sub}(K)$} such that for every $s\leq t\in\mathbb{R}$, we have a nesting of the respective complexes: $F(s)\subseteq F(t)$. [^1] This definition generalizes to any poset as the domain of $F$ rather than $\mathbb{R}$. However, the one-parameter setting is sufficient for our purposes.
A filter on $K$ is a function $f\colon K\to\mathbb{R}$ such that the function $F:\mathbb{R}\to\operatorname{sub}(K)$ defined by
$F(t):=\{\sigma\in K\mid f(\sigma)\leq t\}$ (1)
is a filtration.
Next, we introduce a class of filtrations that are the focus of this work.
@@ 2.1 Lower-Star Filtrations
Let $K$ be a simplicial complex with vertex set $K_{0}$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f:K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function $F:\mathbb{R}\to\operatorname{sub}(K)$, defined as
@@
Let $K$ be a simplicial complex with a map $\ell:K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The vertex height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}:K_{0}\to\mathbb{R}$} defined as $h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$ The directional filtration of $K$ in direction $\mathbf{s}$, denoted as $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
2.2 The Weighted Euler Characteristic Transform
Here, we introduce the main focus of this paper: the WECT.
@@ 2.2 The Weighted Euler Characteristic Transform
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega:K\to\mathbb{R}$ is a function (that we call a weight function).
@@
Originally introduced by Euler in 1758 for convex polyhedra [10], the Euler characteristic (EC) is a topological invariant that readily extends to simplicial complexes. We use an extension of the EC to weighted simplicial complexes: the weighted Euler characteristic.
@@
Note that the Euler characteristic is a special case of the WEC, where the weight function is equal to the constant function $\mathbf{1}:K\to\mathbb{R}$. While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex. To incorporate more information, we apply the WEC to a filtration:
@@
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $F\colon\mathbb{R}\to\operatorname{sub}(K)$ be a filtration of $K$. The weighted Euler characteristic function [^2] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $F$ is the function $\operatorname{WECF}_{F}:\mathbb{R}\to\mathbb{R}$ defined as
@@
When $F$ is induced from a (vertex) filter function $f$, we write $\operatorname{WECF}_{f}:=\operatorname{WECF}_{F}$.
@@
In image analysis tasks (and their three-dimensional analogs), there are two common choices of vertex filters considered. The first (implemented in [13, 24]) uses the pixel intensities as the vertex filter. It then computes the result by applying the (unweighted) Euler characteristic to the corresponding lower-star filtration. We refer to this simply as the ECF of an image. The second, which we refer to as the directional WECF of an image, uses the pixel intensities as weights and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [1, 22, 14].
@@
Figure 1: Filtration of a simplicial complex (1(a))–(1(c)) along direction (1,0), with vertex weights shown in green discs. The weight assigned to a simplex $\sigma$ is the maximum weight of its zero-simplex faces. The corresponding ECF (1(d)) and WECF (1(e)) are displayed in the; see Equation (3). At times 1 and 2, the ECF and WECF differ due to the WECF’s use of the weighted simplices.
Typically, there is no canonical direction vector to choose when computing directional WECFs. Hence, it is common to choose all directions resulting in the weighted Euler characteristic transform (WECT).
@@
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function $\operatorname{WECT}_{\mathbb{K}}:\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$ defined as
@@ 2.3 Existing Frameworks for Computation of the ECF and (W)ECT
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings (resulting in many available software packages). Demeter [1] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT. Next, the SINATRA package [22] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary, unweighted simplicial complexes. Note that as Demeter and SINATRA are not optimized for speed, they are not included in the experimental comparison. Third, EuCalc  [^3] https://github.com/HugoPasse/Eucalc [14] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes. Finally, the DECT package [^4] https://github.com/aidos-lab/dect [17] is a Python package designed to compute the differentiable ECT (DECT) of simplicial complexes. While the DECT is a slightly different variant, the package also includes an implementation for computing the ECT of simplicial or cubical complexes. Although it does not allow for weighted complexes, this package is designed for GPU architectures making it an interesting point of comparison. Furthermore, [18] provides further optimizations for computing the DECT on CUDA GPU architectures, but is consequently a far less general framework.
@@
The FastTopology framework [^5] https://github.com/zavalab/ML/tree/master/FastTopology [13] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF. GPU ECC [24] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems (which is unavailable in our computation environment).
@@
The framework presented in this paper, pyECT, admits several advantages over these existing packages for computing the WECT and ECF. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across axis-aligned cubical complexes, arbitrary cubical complexes, and geometric simplicial complexes; see Table 1. Second, pyECT provides frameworks for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the vectorized operations used in pyECT are highly optimized for modern GPU hardware, which, in most cases, results in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the simplicity of the vectorized approach makes it very advantageous for software implementation, especially in the context of well-maintained, abstracted tensor computation libraries like PyTorch [16, 2]. For example, the core WECT algorithm in the pyECT implementation is only about $25$ lines of code.
@@
One minor disadvantage of pyECT is that it does not compute exact EC values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [12, 23, 15, 1]). While it is possible to compute a vectorization from exact values using EuCalc, it requires an additional computation step that is linear in the size of the discretization.
Table 1: Comparison of acceptable input types of ECT packages.
@@ 2.4 Tensor Operations
Our algorithms almost exclusively use vectorized operations on tensors. These have two main advantages over non-vectorized approaches. First, the vectorized operations we use have highly optimized implementations in PyTorch [16, 2] that run in parallel on both CPUs and GPUs. Second, they support automatic gradient computations, facilitating the use of differentiable versions in our implementations in neural networks. Here, we review tensors and the vectorized operations used in our algorithm.
@@
We exclusively use one-dimensional tensors (i.e., vectors), two-dimensional tensors (i.e., matrices), and three-dimensional tensors. A specific dimension of a tensor is called an axis. We now introduce the tensor operations utilized in our algorithm (and refer the reader to Appendix A for examples).
@@
    Element-wise operations: Any function $f:\mathbb{R}\to\mathbb{R}$ can be applied to tensors element-wise by post-composition. That is, if $T$ is a tensor, then $f(T)$ is the tensor with the same shape as $T$ and values
@@
    In this work, we use the cumulative sum on two-dimensional tensors over the second axis, $m_{1}]]$. However, the cumulative sum can be defined over any axis of a tensor.
@@
    Scatter add: Let $S$ be a tensor with shape $(m_{0},m_{1})$, let $I$ be a tensor with shape $(m_{0},m_{2})$ and values in $m_{1}]]$, and let $T$ be a tensor with shape $(m_{2})$. The function $\texttt{ScatterAdd}(I,T)$ updates $S$ in-place by adding to it, the difference tensor $D$ with shape $(m_{0},m_{1})$ and values

### v3

# Tensor Computation of Euler Characteristic Functions and Transforms

@@ Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are either not optimized for GPU computation or do not scale to higher-dimensional settings. In this work, we present a tensor-based framework for computing such topological descriptors which is highly optimized for GPU architectures and works in full generality across simplicial and cubical complexes of arbitrary dimension. Experimentally, the framework demonstrates significant speedups over existing methods when computing the WECT and ECF across a variety of two- and three-dimensional datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT. [^1] https://github.com/compTAG/pyECT
@@ 1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant, the Euler characteristic [10]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [9] and Euler characteristic functions [21].
Topological transforms [24] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a directional filtration of $K$ is constructed and then a topological invariant is applied to that filtration. Applying the Euler characteristic (EC) to a single directional filtration results in an Euler characteristic function (ECF), while applying it to each directional filtration in $\mathbb{S}^{n-1}$ results in the Euler characteristic transform (ECT) [24, 15]. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [11, 19]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
Useful variants of the ECF and ECT are the weighted Euler characteristic function (WECF) and the weighted Euler characteristic transform (WECT) for weighted complexes [12]. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [12]. Moreover, the WECT has been employed in a broad range of applications, including: in oncology for predicting survival rates of patients with glioblastoma multiforme brain tumors [6], in plant biology for classifying seed phenotypes [1], and in image classification tasks [3, 4, 12]. Numerous software packages already exist for computing the WECT and its variants, including those presented in [1, 6, 13, 14, 17, 18]. While many of these packages provide useful tools, they are either not optimized for GPU architectures or are limited in the types (or dimension) of complexes they can process.
In this work, we provide a fully tensor-based framework for computing WECFs and WECTs. Our implementation applies to arbitrary geometric simplicial and cubical complexes in Euclidean spaces. The tensor operations we use are highly optimized for fast computation on modern GPU hardware, providing significant speedups over existing implementations.
@@ 2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to (abstract) simplicial complexes; however, our results and code also work for cubical complexes.
For a simplicial complex $K$ and for any $i\in\{0,1,\ldots,\dim(K)\}$, we denote the set of $i$-simplices of $K$ by $K_{i}$. We use the notation $\operatorname{sub}(K)$ for the set of subcomplexes of $K$. Furthermore, we define $k=|K|$ and $k_{i}=|K_{i}|$.
@@
Let $K$ be a simplicial complex with vertex set $K_{0}$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f\colon K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function $F\colon\mathbb{R}\to\operatorname{sub}(K)$, defined as
@@
Let $K$ be a simplicial complex with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The vertex height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}\colon K_{0}\to\mathbb{R}$} defined as $h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$ The directional filtration of $K$ in direction $\mathbf{s}$, denoted as $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
2.1 Weighted Euler Characteristic Functions and Transforms
Here, we introduce weighted Euler characteristic functions and transforms. For a more thorough review, see the survey paper [15].
@@ 2.1 Weighted Euler Characteristic Functions and Transforms
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega\colon K\to\mathbb{R}$ is a function called a weight function.
@@
Note that the Euler characteristic is a special case of the WEC, where the weight function is equal to the constant function $\mathbf{1}\colon K\to\mathbb{R}$. While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex. To incorporate more information, we apply the WEC to a filtration:
@@
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $f\colon K_{0}\to\mathbb{R}$ be a vertex filter of $K$ with lower-star filtration $F$. The weighted Euler characteristic function [^2] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $f$ is the function $\operatorname{WECF}_{f}\colon\mathbb{R}\to\mathbb{R}$ defined as
@@
In image analysis tasks (and their three-dimensional analogs), there are two common (W)ECFs considered. The first (implemented in [13, 26, 25]) uses the pixel intensities as the vertex filter. We refer to the resulting ECF simply as the ECF of an image. The second, which we refer to as the directional WECF of an image, uses the pixel intensities as weights and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [1, 23, 14].
@@
Figure 1: Filtration of a simplicial complex (1(a))–(1(c)) along direction (1,0), with vertex weights shown in green discs. The weight assigned to a simplex is the maximum weight of its vertices. The corresponding ECF and WECF are displayed in (1(d)) and (1(e)) respectively.
Typically, there is no canonical direction vector to choose when computing directional WECFs. Hence, it is common to choose all directions, resulting in the weighted Euler characteristic transform (WECT).
@@
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function
$\operatorname{WECT}_{\mathbb{K}}\colon\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$
defined as
@@ 2.2 Existing Frameworks for Computation of the ECF and (W)ECT
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings (resulting in many available software packages). Demeter [1] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT. Next, the SINATRA package [23] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary, unweighted simplicial complexes. Note that as Demeter and SINATRA are not optimized for speed, they are not included in the experimental comparison. Third, EuCalc  [^3] https://github.com/HugoPasse/Eucalc [14] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes. Finally, the DECT package [^4] https://github.com/aidos-lab/dect [17] is a Python package designed to compute the differentiable ECT (DECT) of simplicial complexes. While the DECT is a slightly different variant, the package also includes an implementation for computing the ECT of simplicial or cubical complexes. Although it does not allow for weighted complexes, this package is designed for GPU architectures, making it an interesting point of comparison. Furthermore, [18] provides further optimizations for computing the DECT on CUDA GPUs, but is consequently a far less general framework.
@@
The FastTopology framework [^5] https://github.com/zavalab/ML/tree/master/FastTopology [13] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF. GPU ECC [26, 25] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems (which is unavailable in our computation environment).
@@
The framework presented in this paper, pyECT, admits several advantages over these existing packages for computing the WECT and ECF. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across arbitrary weighted simplicial and cubical complexes; see Table 1. Second, pyECT provides methods for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the tensor operations used in pyECT are highly optimized for modern GPU hardware, resulting in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the tensor-based approach used in pyECT yields simple code (our core WECT algorithm is around $25$ lines of code) that is easily readable, maintable, and customizable.
@@
One minor disadvantage of pyECT is that it does not compute exact EC values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [12, 24, 15, 1]).
Table 1: Comparison of supported input types for ECT packages.
@@ 2.3 Tensor Operations
Our algorithm almost exclusively uses tensor operations. These operations have highly optimized implementations in PyTorch [16, 2] that run in parallel on both CPUs and GPUs. Here, we review tensors and the tensor operations used in our algorithm (see Appendix A for examples).
@@
We exclusively use one-dimensional tensors (vectors), two-dimensional tensors (matrices), and three-dimensional tensors. A specific dimension of a tensor is called an axis. We now introduce the tensor operations utilized in our algorithm.
@@
    Element-wise operations: Any function $f\colon\mathbb{R}\to\mathbb{R}$ can be applied to tensors element-wise by post-composition. That is, if $T$ is a tensor, then $f(T)$ is the tensor with the same shape as $T$ and values
@@
    In this work, we use the cumulative sum on two-dimensional tensors along the second axis. However, the cumulative sum can be defined over any axis of a tensor.
@@
    Scatter add: Let $S$ be a tensor with shape $(m_{0},m_{1})$, let $I$ be a tensor with shape $(m_{0},m_{2})$ and values in $m_{1}]]$, and let $T$ be a tensor with shape $(m_{2})$. The function $\texttt{ScatterAdd}(I,T)$ updates $S$ in-place by adding to it the difference tensor $D$ with shape $(m_{0},m_{1})$ and values